# Concentration and fluctuations of sine–Gordon measure around topological multi-soliton manifold

### v1

@@ 9.1. Asymptotic behavior of the covariance
$v(x)=\mathbf{P}(x)\langle v,g\rangle+w(x),$ (9.3)
@@
where the variance $\sigma_{g}^{2}$ is given by
$\sigma_{g}^{2}=\mathbb{E}_{\nu^{\perp}_{\xi_{1},\dots,\xi_{k}}}\big[|\langle v,g\rangle|^{2}\big].$
Before presenting the proof of Proposition 9.1, we study the following lemma, which provides control on the variance $\sigma_{g}^{2}$.
@@
$\mathbb{E}_{\nu^{\perp}_{\xi_{1},\dots,\xi_{k}}}\big[|\langle v,g\rangle|^{2}\big]=\langle g,G^{D}_{\xi_{1},\dots,\xi_{k}}g\rangle,$ (9.4)
where $G^{D}_{\xi_{1},\dots,\xi_{k}}$ is the Green’s function for the covariance operator
$C_{\xi_{1},\dots,\xi_{k}}=\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}}(-\partial_{x}^{2}+1+W)^{-1}\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}},$
defined in (5.11). Here
$W:=-2\sum_{j=1}^{k}\text{sech}^{2}(\cdot-\xi_{j})$
is the perturbation term, localized near each soliton center. Using the resolvent identity, we interpret the Schrödinger operator $C_{\xi_{1},\dots,\xi_{k}}$ as a perturbation of the projected Ornstein–Uhlenbeck operator $\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}}(-\partial_{x}^{2}+1)^{-1}\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}}$ as follows
$G^{D}_{\xi_{1},\dots,\xi_{k}}(x,y)=G_{\text{OU},\xi_{1},\dots,\xi_{k}}^{D}(x,y)+(G_{\text{OU},\xi_{1},\dots,\xi_{k}}^{D}WG^{D}_{\xi_{1},\dots,\xi_{k}})(x,y)$ (9.5)
where
$G_{\text{OU},\xi_{1},\dots,\xi_{k}}^{D}(x,y)=(\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}}(-\partial_{x}^{2}+1)^{-1}\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}})(x,y).$
We first study the leading order term $G_{\text{OU},\xi_{1},\dots,\xi_{k}}^{D}$ in (9.5). By expanding the projection operator $\mathbf{P}_{V_{\xi_{1}\,\dots,\xi_{k}}}$, defined in (5.12), we write
$\hphantom{X}-\sum_{i=1}^{k}\mathbf{P}_{i}(-\partial_{x}^{2}+1)^{-1}-\sum_{j=1}^{k}(-\partial_{x}^{2}+1)^{-1}\mathbf{P}_{j},$ (9.6)
where $\mathbf{P}_{j}=\langle\cdot,t_{j}\rangle t_{j}$ is the projection onto the tangential direction. We begin with the first projected term $\mathbf{P}_{i}(-\partial_{x}^{2}+1)^{-1}\mathbf{P}_{j}$. Note that
@@
$|\langle g,t_{j}\rangle|\lesssim\bigg|\int_{\operatorname*{supp}g}g(x)e^{-|x-\xi_{j}|}dx\bigg|\lesssim\|g\|_{L^{2}}e^{-\textup{dist}(\operatorname*{supp}g,\xi_{j})}\lesssim e^{-\varepsilon^{-\frac{1}{2}+2\eta}}.$ (9.7)
@@
$|\langle g,\mathbf{P}_{i}(-\partial_{x}^{2}+1)^{-1}\mathbf{P}_{j}g\rangle|\lesssim e^{-\varepsilon^{-\frac{1}{2}+2\eta}}.$ (9.8)
@@
$|\langle g,\mathbf{P}_{i}(-\partial_{x}^{2}+1)^{-1}g\rangle|\lesssim e^{-\varepsilon^{-\frac{1}{2}+2\eta}}\quad\textup{and}\quad|\langle g,(-\partial_{x}^{2}+1)^{-1}\mathbf{P}_{j}g\rangle|\lesssim e^{-\varepsilon^{-\frac{1}{2}+2\eta}}.$ (9.9)
@@
$\langle g,G^{D}_{\textup{OU},\xi_{1},\dots,\xi_{k}}g\rangle=\langle g,(-\partial_{x}^{2}+1)^{-1}g\rangle+O(e^{-\varepsilon^{-\frac{1}{2}+2\eta}}).$ (9.10)
@@
We now study the perturbation term $G_{\text{OU},\xi_{1},\dots,\xi_{k}}^{D}WG^{D}_{\xi_{1},\dots,\xi_{k}}$ in (9.5), which is also an error term. We expand
@@
$\lesssim\sum_{j=1}^{k}\bigg|\int_{-L_{\varepsilon}}^{L_{\varepsilon}}g(x)\bigg(\int_{-L_{\varepsilon}}^{L_{\varepsilon}}G_{\text{OU},\xi_{1},\dots,\xi_{k}}^{D}(x,z)e^{-|z-\xi_{j}|}\bigg(\int_{-L_{\varepsilon}}^{L_{\varepsilon}}G^{D}_{\xi_{1},\dots,\xi_{k}}(z,y)g(y)dy\bigg)dz\bigg)dx\bigg|,$ (9.11)
where we used $\textup{sech}^{2}(z-\xi_{j})\lesssim e^{-|z-\xi_{j}|}$. To use the correlation decay, we separate the Green’s function into its bulk part and boundary parts, as follows, based on Lemma 8.2 and Proposition 8.3,
$G_{\text{OU},\xi_{1},\dots,\xi_{k}}^{D}(x,z)$ $=ce^{-|x-z|}+O(e^{-2(L_{\varepsilon}-\max\{|x|,|z|\})})$ (9.12)
$G^{D}_{\xi_{1},\dots,\xi_{k}}(z,y)$ $=ce^{-|z-y|}+O(e^{-2(L_{\varepsilon}-\max\{|z|,|y|\})}).$ (9.13)
We first focus on the bulk effect $e^{-|x-z|}$ and $e^{-|z-y|}$ by plugging them into (9.11)
@@
$\lesssim\sum_{j=1}^{k}L_{\varepsilon}e^{-\frac{1}{2}\text{dist}(\operatorname*{supp}g,\xi_{j})}\|g\|_{L^{1}}\|g\|_{L^{1}}$
$\lesssim L_{\varepsilon}e^{-\frac{1}{2}\varepsilon^{-\frac{1}{2}+2\eta}}\lesssim e^{-\frac{1}{4}\varepsilon^{-\frac{1}{2}+2\eta}},$ (9.14)
where in the last step, we use the condition $\textup{dist}(\operatorname*{supp}g,\xi_{j})\geq\varepsilon^{-\frac{1}{2}+2\eta}$ for every $1\leq j\leq k$.
For the boundary effects $e^{-2(L_{\varepsilon}-\max\{|x|,|z|\})}$ and $e^{-2(L_{\varepsilon}-\max\{|z|,|y|\})}$ in (9.12) and (9.13), following the arguments in (8.10) and (8.12), we obtain the error term $e^{-c\varepsilon^{-\frac{1}{2}+2\eta}}$. Combining (9.11), (9.14), and the boundary error $e^{-c\varepsilon^{-\frac{1}{2}+2\eta}}$ yields
$|\langle g,(G_{\text{OU},\xi_{1},\dots,\xi_{k}}^{D}WG^{D}_{\xi_{1},\dots,\xi_{k}})g\rangle|\lesssim e^{-\frac{1}{4}\varepsilon^{-\frac{1}{2}+2\eta}}+e^{-c\varepsilon^{-\frac{1}{2}+2\eta}}$ (9.15)
@@ 9.2. Ornstein–Uhlenbeck limit on the normal space
$=\hskip 1.42262pt\text{I}\hskip 1.42262pt_{1}+\hskip 1.42262pt\text{I}\hskip 1.42262pt_{2},$ (9.16)
@@
$\Big|\frac{1}{\varepsilon}\mathcal{E}_{\xi_{1},\dots,\xi_{k}}(\sqrt{\varepsilon}v)\Big|\leq\frac{1}{\varepsilon}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}|\sqrt{\varepsilon}v|^{3}dx$ $\leq\sqrt{\varepsilon}L_{\varepsilon}\|v\|_{L^{\infty}}^{3}$
$\lesssim\varepsilon^{0+}|\log\varepsilon|^{\frac{3}{2}}=\varepsilon^{0+},$
@@
$\textup{Det}_{\xi_{1},\dots,\xi_{k}}(\sqrt{\varepsilon}v)=1+O(\|\sqrt{\varepsilon}v\|_{L^{\infty}}^{k})=1+O(\varepsilon^{\frac{k}{2}-})$ (9.17)
@@
$\hskip 1.42262pt\text{I}\hskip 1.42262pt_{2}=O(\varepsilon^{0+}),$ (9.18)
uniformly in $\xi_{1},\dots,\xi_{k}\in[-\overline{L}_{\varepsilon},\overline{L}_{\varepsilon}]^{k}$.
Regarding the main term $\hskip 1.42262pt\text{I}\hskip 1.42262pt_{1}$, we use (9.17) and the tail probability estimates in Lemmas 8.5 and (6.28) to obtain
$\hskip 1.42262pt\text{I}\hskip 1.42262pt_{1}=\mathbb{E}_{\nu^{\perp}_{\xi_{1},\dots,\xi_{k}}}\Big[e^{i\langle v,g\rangle}\Big]+O(\varepsilon^{\frac{k}{2}-})+O(e^{-cK^{2}|\log L_{\varepsilon}|})+O(e^{-\frac{c\delta_{\varepsilon}^{2}}{\varepsilon}}),$ (9.19)
where $\delta_{\varepsilon}=\sqrt{\varepsilon\log\frac{1}{\varepsilon}}$. Using the orthogonal decomposition $v=\mathbf{P}(x)\langle v,g\rangle+w$ in (9.3), where $\langle v,g\rangle$ and $w$ are independent Gaussian, we write
@@
$=\exp\Big\{-\frac{1}{2}\langle g,(-\partial_{x}^{2}+1)^{-1}g\rangle\cdot\big(1+O(e^{-c\varepsilon^{-\frac{1}{2}+2\eta}})\big)\Big\},$ (9.20)
where we used Lemma 9.3
$\sigma_{g}^{2}=\mathbb{E}_{\nu^{\perp}_{m_{\xi_{1},\dots,\xi_{k}}}}\big[|\langle v,g\rangle|^{2}\big]=\langle g,(-\partial_{x}^{2}+1)^{-1}g\rangle\cdot\big(1+O(e^{-c\varepsilon^{-\frac{1}{2}+2\eta}})\big).$
By combining (9.16), (9.18), (9.19), and (9.20), we obtain
@@
$\mathbb{E}_{\nu^{\perp}_{\xi_{1},\dots,\xi_{k}}}\Big[e^{\frac{1}{\varepsilon}\mathcal{E}_{\xi_{1},\dots,\xi_{k}}(\sqrt{\varepsilon}v)}\textup{Det}_{\xi_{1},\dots,\xi_{k}}(\sqrt{\varepsilon}v),\;A_{\varepsilon},\;\|\sqrt{\varepsilon}v\|_{L^{2}}<\delta_{\varepsilon}\Big]=1+O(\varepsilon^{0+}),$ (9.21)
uniformly in $\xi_{1},\dots,\xi_{k}\in[-\overline{L}_{\varepsilon},\overline{L}_{\varepsilon}]^{k}$, where $\delta_{\varepsilon}=\sqrt{\varepsilon\log\frac{1}{\varepsilon}}$. Note that in (9.21) we do not need the conditions $\textup{dist}(\operatorname*{supp}g,\xi_{j})\geq\varepsilon^{-\frac{1}{2}+2\eta}$ for $j=1,\dots k$. These conditions are only used in (9.20), where we used Lemma 9.3. Note that if $e^{it}$ is replaced by $1$, then (9.20) is immediately equal to $1$.
@@
where $T^{\varepsilon}(\phi)=\varepsilon^{-\frac{1}{2}}(\phi-\pi^{\varepsilon}(\phi))$ and $\pi^{\varepsilon}$ denotes the projection onto the multi-soliton manifold $\mathcal{M}^{\varepsilon,\geq(1+\eta)d_{\varepsilon}}_{Q}$, defined in (4.15). Using Lemma 7.1, we decompose the integral into the large-deviation events and the main term $\mathcal{J}_{1}$ as follows
$\int e^{i\langle T^{\varepsilon}(\phi),g\rangle}\rho_{\varepsilon}^{Q}(d\phi)=\mathcal{J}_{1}+\textup{O}(e^{-c\log\frac{1}{\varepsilon}})$ (10.1)
@@
$\mathcal{J}_{1}$ $=\int_{\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon,\,\geq(1+\eta)d_{\varepsilon}})<\delta_{\varepsilon}\}}e^{\langle T^{\varepsilon}(\phi),g\rangle}\rho_{\varepsilon}^{Q}(d\phi),$
Here $\delta_{\varepsilon}=\eta\sqrt{\varepsilon\log\frac{1}{\varepsilon}}$ and $d_{\varepsilon}=\big|\log\sqrt{\varepsilon\log\frac{1}{\varepsilon}}\big|$, as defined in (6.28) and (6.29). By using the coordinate expression $\phi=m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}}+\sqrt{\varepsilon}v$ and Proposition 7.4, we write
$\mathcal{J}_{1}=\overline{Z}_{\varepsilon}^{-1}\mathop{\int\cdots\int}\limits_{-\overline{L}_{\varepsilon}\leq\xi_{1}\leq\cdots\leq\xi_{Q}\leq\overline{L}_{\varepsilon}}\mathcal{F}_{\xi_{1},\dots,\xi_{Q}}(e^{i\langle v,g\rangle})\mathbf{1}_{\{\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq(1+\eta)d_{\varepsilon}\}}d\xi_{1}\dots d\xi_{Q},$ (10.2)
where
@@
$\overline{Z}_{\varepsilon}=\mathop{\int\cdots\int}\limits_{-\overline{L}_{\varepsilon}\leq\xi_{1}\leq\cdots\leq\xi_{Q}\leq\overline{L}_{\varepsilon}}\mathcal{F}_{\xi_{1},\dots,\xi_{Q}}(1)\mathbf{1}_{\{\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq(1+\eta)d_{\varepsilon}\}}d\xi_{1}\dots,d\xi_{Q}.$ (10.3)
We now split the main term $J_{1}$ in (10.2) into two pats, $\Xi_{\varepsilon}$ and $\Xi_{\varepsilon}^{c}$ as follows
$\mathcal{J}_{1}$ $=\overline{Z}_{\varepsilon}^{-1}\int\cdots\int_{\Xi_{\varepsilon}}\mathcal{F}_{\xi_{1},\dots,\xi_{Q}}(e^{i\langle v,g\rangle})\mathbf{1}_{\{\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq(1+\eta)d_{\varepsilon}\}}d\xi_{1}\dots,d\xi_{Q}$
$\hphantom{X}+\overline{Z}_{\varepsilon}^{-1}\int\cdots\int_{\Xi_{\varepsilon}^{c}}\mathcal{F}_{\xi_{1},\dots,\xi_{Q}}(e^{i\langle v,g\rangle})\mathbf{1}_{\{\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq(1+\eta)d_{\varepsilon}\}}d\xi_{1}\dots,d\xi_{Q}$
$=\mathcal{J}_{1}^{(1)}+\mathcal{J}_{1}^{(2)},$ (10.4)
@@
$\Xi_{\varepsilon}=\big\{-\overline{L}_{\varepsilon}\leq\xi_{1}\leq\cdots\leq\xi_{Q}\leq\overline{L}_{\varepsilon}:\textup{dist}(\operatorname*{supp}g,\xi_{j})\geq\varepsilon^{-\frac{1}{2}+2\eta}\;\;\textup{for every $j$}\big\}.$ (10.5)
@@
$\mathcal{F}_{\xi_{1},\dots,\xi_{Q}}(e^{i\langle v,g\rangle})=\exp\Big\{-\frac{1}{2}\langle g,(-\partial_{x}^{2}+1)^{-1}g\rangle_{L^{2}(\mathbb{R})}\Big\}\cdot(1+O(\varepsilon^{0+})),$ (10.6)
@@
$\mathcal{F}_{\xi_{1},\dots,\xi_{Q}}(1)=1+O(\varepsilon^{0+}),$ (10.7)
uniformly in $(\xi_{1},\dots,\xi_{Q})\in[-\overline{L}_{\varepsilon},\overline{L}_{\varepsilon}]^{k}$.
Combining (10.4), (10.6), (10.3), and (10.7) yields that
$\mathcal{J}_{1}^{(1)}$ $=e^{-\frac{1}{2}\|g\|^{2}_{H^{-1}(\mathbb{R})}}\cdot\frac{\big|\Xi_{\varepsilon}\cap\{\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq(1+\eta)d_{\varepsilon}\}\big|}{\big|\big\{-\overline{L}_{\varepsilon}\leq\xi_{1}\leq\cdots\leq\xi_{Q}\leq\overline{L}_{\varepsilon}\big\}\cap\{\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq(1+\eta)d_{\varepsilon}\}\big|}(1+O(\varepsilon^{0+})).$ (10.8)

### v2

@@ 9.1. Asymptotic behavior of the covariance
(9.3) $v(x)=\mathbf{P}(x)\langle v,g\rangle+w(x),$
@@
where $\sigma_{g}^{2}=\mathbb{E}_{\nu^{\perp}_{\xi_{1},\dots,\xi_{k}}}\big[|\langle v,g\rangle|^{2}\big]$. Before presenting the proof of Proposition 9.1, we study the following lemma, which provides control on the variance $\sigma_{g}^{2}$.
@@
(9.4) $\mathbb{E}_{\nu^{\perp}_{\xi_{1},\dots,\xi_{k}}}\big[|\langle v,g\rangle|^{2}\big]=\langle g,G^{D}_{\xi_{1},\dots,\xi_{k}}g\rangle,$
where $G^{D}_{\xi_{1},\dots,\xi_{k}}$ is the Green’s function, defined in (8.2). Recall the resolvent identity (8.17) for $G^{D}_{\xi_{1},\dots,\xi_{k}}$. We first study the leading order term $G_{\text{OU},\xi_{1},\dots,\xi_{k}}^{D}$ in (8.17). By expanding the projection $\mathbf{P}_{V_{\xi_{1}\,\dots,\xi_{k}}}$ as in (8.4), we begin with the first projected term $\mathbf{P}_{i}(-\partial_{x}^{2}+1)^{-1}\mathbf{P}_{j}$. Note that
@@
(9.5) $|\langle g,t_{j}\rangle|\lesssim\bigg|\int_{\operatorname*{supp}g}g(x)e^{-|x-\xi_{j}|}dx\bigg|\lesssim\|g\|_{L^{2}}e^{-\textup{dist}(\operatorname*{supp}g,\xi_{j})}\lesssim e^{-\varepsilon^{-\frac{1}{2}+2\eta}}.$
@@
(9.6) $|\langle g,\mathbf{P}_{i}(-\partial_{x}^{2}+1)^{-1}\mathbf{P}_{j}g\rangle|\lesssim e^{-\varepsilon^{-\frac{1}{2}+2\eta}}.$
@@
(9.7) $|\langle g,\mathbf{P}_{i}(-\partial_{x}^{2}+1)^{-1}g\rangle|\lesssim e^{-\varepsilon^{-\frac{1}{2}+2\eta}}\quad\textup{and}\quad|\langle g,(-\partial_{x}^{2}+1)^{-1}\mathbf{P}_{j}g\rangle|\lesssim e^{-\varepsilon^{-\frac{1}{2}+2\eta}}.$
@@
(9.8) $\langle g,G^{D}_{\textup{OU},\xi_{1},\dots,\xi_{k}}g\rangle=\langle g,(-\partial_{x}^{2}+1)^{-1}g\rangle+O(e^{-\varepsilon^{-\frac{1}{2}+2\eta}}).$
@@
We now study the perturbation term $G_{\text{OU},\xi_{1},\dots,\xi_{k}}^{D}WG^{D}_{\xi_{1},\dots,\xi_{k}}$ in (8.17). We expand
@@
(9.9) $\lesssim\sum_{j=1}^{k}\bigg|\int_{-L_{\varepsilon}}^{L_{\varepsilon}}g(x)\bigg(\int_{-L_{\varepsilon}}^{L_{\varepsilon}}G_{\text{OU},\xi_{1},\dots,\xi_{k}}^{D}(x,z)e^{-|z-\xi_{j}|}\bigg(\int_{-L_{\varepsilon}}^{L_{\varepsilon}}G^{D}_{\xi_{1},\dots,\xi_{k}}(z,y)g(y)dy\bigg)dz\bigg)dx\bigg|,$
where we used $\textup{sech}^{2}(z-\xi_{j})\lesssim e^{-|z-\xi_{j}|}$. To use the correlation decay, we separate $G_{\text{OU},\xi_{1},\dots,\xi_{k}}^{D}(x,z)$ and $G^{D}_{\xi_{1},\dots,\xi_{k}}(z,y)$ into its bulk and boundary parts, based on Lemma 8.2 and Proposition 8.3 We first focus on the bulk effect $e^{-|x-z|}$ and $e^{-|z-y|}$ by plugging them into (9.9)
@@
(9.10) $\lesssim\sum_{j=1}^{k}L_{\varepsilon}e^{-\frac{1}{2}\text{dist}(\operatorname*{supp}g,\xi_{j})}\|g\|_{L^{1}}\|g\|_{L^{1}}\lesssim L_{\varepsilon}e^{-\frac{1}{2}\varepsilon^{-\frac{1}{2}+2\eta}}\lesssim e^{-\frac{1}{4}\varepsilon^{-\frac{1}{2}+2\eta}},$
where in the last step, we use the condition $\textup{dist}(\operatorname*{supp}g,\xi_{j})\geq\varepsilon^{-\frac{1}{2}+2\eta}$ for every $1\leq j\leq k$. For the boundary effects $e^{-2(L_{\varepsilon}-\max\{|x|,|z|\})}$ and $e^{-2(L_{\varepsilon}-\max\{|z|,|y|\})}$ in Lemma 8.2 and Proposition 8.3, following the arguments in (8.10) and (8.12), we obtain the error term $e^{-c\varepsilon^{-\frac{1}{2}+2\eta}}$. Combining (9.9), (9.10), and the boundary error $e^{-c\varepsilon^{-\frac{1}{2}+2\eta}}$ yields
(9.11) $|\langle g,(G_{\text{OU},\xi_{1},\dots,\xi_{k}}^{D}WG^{D}_{\xi_{1},\dots,\xi_{k}})g\rangle|\lesssim e^{-\frac{1}{4}\varepsilon^{-\frac{1}{2}+2\eta}}+e^{-c\varepsilon^{-\frac{1}{2}+2\eta}}$
@@ 9.2. Ornstein–Uhlenbeck limit on the normal space
(9.12) $=\hskip 1.42262pt\text{I}\hskip 1.42262pt_{1}+\hskip 1.42262pt\text{I}\hskip 1.42262pt_{2},$
@@
$\Big|\frac{1}{\varepsilon}\mathcal{E}_{\xi_{1},\dots,\xi_{k}}(\sqrt{\varepsilon}v)\Big|\leq\frac{1}{\varepsilon}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}|\sqrt{\varepsilon}v|^{3}dx$ $\leq\sqrt{\varepsilon}L_{\varepsilon}\|v\|_{L^{\infty}}^{3}\lesssim\varepsilon^{0+}|\log\varepsilon|^{\frac{3}{2}}=\varepsilon^{0+},$
@@
(9.13) $\textup{Det}_{\xi_{1},\dots,\xi_{k}}(\sqrt{\varepsilon}v)=1+O(\|\sqrt{\varepsilon}v\|_{L^{\infty}}^{k})=1+O(\varepsilon^{\frac{k}{2}-})$
@@
(9.14) $\hskip 1.42262pt\text{I}\hskip 1.42262pt_{2}=O(\varepsilon^{0+}),$
uniformly in $\xi_{1},\dots,\xi_{k}\in[-\overline{L}_{\varepsilon},\overline{L}_{\varepsilon}]^{k}$. Regarding the main term $\hskip 1.42262pt\text{I}\hskip 1.42262pt_{1}$, we use (9.13) and the tail probability estimates in Lemmas 8.5 and (6.29) to obtain
(9.15) $\hskip 1.42262pt\text{I}\hskip 1.42262pt_{1}=\mathbb{E}_{\nu^{\perp}_{\xi_{1},\dots,\xi_{k}}}\Big[e^{i\langle v,g\rangle}\Big]+O(\varepsilon^{\frac{k}{2}-})+O(e^{-cK^{2}|\log L_{\varepsilon}|})+O(e^{-\frac{c\delta_{\varepsilon}^{2}}{\varepsilon}}),$
where $\delta_{\varepsilon}=\eta\sqrt{\varepsilon\log\frac{1}{\varepsilon}}$. Using the orthogonal decomposition $v=\mathbf{P}(x)\langle v,g\rangle+w$ in (9.3), where $\langle v,g\rangle$ and $w$ are independent Gaussian, we write
@@
(9.16) $=\exp\Big\{-\frac{1}{2}\langle g,(-\partial_{x}^{2}+1)^{-1}g\rangle\cdot\big(1+O(e^{-c\varepsilon^{-\frac{1}{2}+2\eta}})\big)\Big\},$
where we used Lemma 9.3. By combining (9.12), (9.14), (9.15), and (9.16), we obtain
@@
(9.17) $\mathbb{E}_{\nu^{\perp}_{\xi_{1},\dots,\xi_{k}}}\Big[e^{\frac{1}{\varepsilon}\mathcal{E}_{\xi_{1},\dots,\xi_{k}}(\sqrt{\varepsilon}v)}\textup{Det}_{\xi_{1},\dots,\xi_{k}}(\sqrt{\varepsilon}v),\;A_{\varepsilon},\;\|\sqrt{\varepsilon}v\|_{L^{2}}<\delta_{\varepsilon}\Big]=1+O(\varepsilon^{0+}),$
uniformly in $\xi_{1},\dots,\xi_{k}\in[-\overline{L}_{\varepsilon},\overline{L}_{\varepsilon}]^{k}$, where $\delta_{\varepsilon}=\eta\sqrt{\varepsilon\log\frac{1}{\varepsilon}}$. Note that in (9.17) we do not need the conditions $\textup{dist}(\operatorname*{supp}g,\xi_{j})\geq\varepsilon^{-\frac{1}{2}+2\eta}$ for $j=1,\dots k$. These conditions are only used in (9.16), where we used Lemma 9.3. Note that if $e^{it}$ is replaced by $1$, then (9.16) is immediately equal to $1$.
@@
where $T^{\varepsilon}(\phi)=\varepsilon^{-\frac{1}{2}}(\phi-\pi^{\varepsilon}(\phi))$ and $\pi^{\varepsilon}$ denotes the projection onto the multi-soliton manifold $\mathcal{M}^{\varepsilon,\geq d_{\varepsilon}}_{Q}$, defined in (4.15). Using Lemma 7.1, we decompose the integral into the large-deviation events and the main term $\mathcal{J}_{1}$ as follows
(10.1) $\int e^{i\langle T^{\varepsilon}(\phi),g\rangle}\rho_{\varepsilon}^{Q}(d\phi)=\mathcal{J}_{1}+\textup{O}(e^{-c\log\frac{1}{\varepsilon}})$
@@
$\mathcal{J}_{1}$ $=\int_{\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon,\,\geq d_{\varepsilon}})<\delta_{\varepsilon}\}}e^{\langle T^{\varepsilon}(\phi),g\rangle}\rho_{\varepsilon}^{Q}(d\phi).$
Here $\delta_{\varepsilon}=\eta\sqrt{\varepsilon\log\frac{1}{\varepsilon}}$ and $d_{\varepsilon}=\big|\log(\varepsilon\log\frac{1}{\varepsilon})\big|$, as defined in (6.29) and (6.30). By using the coordinate expression $\phi=m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}}+\sqrt{\varepsilon}v$ and Proposition 7.2, we write
(10.2) $\mathcal{J}_{1}=\overline{Z}_{\varepsilon}^{-1}\int\cdots\int_{\Delta_{Q}}\mathcal{F}_{\xi_{1},\dots,\xi_{Q}}(e^{i\langle v,g\rangle})\mathbf{1}_{\{\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq d_{\varepsilon}\}}d\xi_{1}\dots d\xi_{Q},$
where $\Delta_{Q}=\{-\overline{L}_{\varepsilon}\leq\xi_{1}\leq\cdots\leq\xi_{Q}\leq\overline{L}_{\varepsilon}\}$ is defined as in (4.17) and
@@
(10.3) $\overline{Z}_{\varepsilon}=\int\cdots\int_{\Delta_{Q}}\mathcal{F}_{\xi_{1},\dots,\xi_{Q}}(1)\mathbf{1}_{\{\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq d_{\varepsilon}\}}d\xi_{1}\dots,d\xi_{Q}.$
We now split the main term $J_{1}$ in (10.2) into two parts, $\Xi_{\varepsilon}$ and $\Xi_{\varepsilon}^{c}$ as follows
$\mathcal{J}_{1}$ $=\overline{Z}_{\varepsilon}^{-1}\int\cdots\int_{\Xi_{\varepsilon}}\mathcal{F}_{\xi_{1},\dots,\xi_{Q}}(e^{i\langle v,g\rangle})\mathbf{1}_{\{\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq d_{\varepsilon}\}}d\xi_{1}\dots,d\xi_{Q}$
$\hphantom{X}+\overline{Z}_{\varepsilon}^{-1}\int\cdots\int_{\Xi_{\varepsilon}^{c}}\mathcal{F}_{\xi_{1},\dots,\xi_{Q}}(e^{i\langle v,g\rangle})\mathbf{1}_{\{\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq d_{\varepsilon}\}}d\xi_{1}\dots,d\xi_{Q}$
(10.4) $=\mathcal{J}_{1}^{(1)}+\mathcal{J}_{1}^{(2)},$
@@
(10.5) $\Xi_{\varepsilon}=\big\{(\xi_{1},\dots,\xi_{Q})\in\Delta_{Q}:\textup{dist}(\operatorname*{supp}g,\xi_{j})\geq\varepsilon^{-\frac{1}{2}+2\eta}\;\;\textup{for every $j$}\big\}.$
@@
(10.6) $\mathcal{F}_{\xi_{1},\dots,\xi_{Q}}(e^{i\langle v,g\rangle})=\exp\Big\{-\frac{1}{2}\langle g,(-\partial_{x}^{2}+1)^{-1}g\rangle_{L^{2}(\mathbb{R})}\Big\}\cdot(1+O(\varepsilon^{0+})),$
@@
(10.7) $\mathcal{F}_{\xi_{1},\dots,\xi_{Q}}(1)=1+O(\varepsilon^{0+}),$
uniformly in $(\xi_{1},\dots,\xi_{Q})\in\Delta_{Q}$. Combining (10.4), (10.6), (10.3), and (10.7) yields that
(10.8) $\mathcal{J}_{1}^{(1)}$ $=e^{-\frac{1}{2}\|g\|^{2}_{H^{-1}(\mathbb{R})}}\cdot\frac{\big|\Xi_{\varepsilon}\cap\{\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq d_{\varepsilon}\}\big|}{\big|\Delta_{Q}\cap\{\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq d_{\varepsilon}\}\big|}(1+O(\varepsilon^{0+})).$